camrynh0801:

A triangle with vertices at A(0, 0), B(0, 4), and C(6, 0) is dilated to yield a triangle with vertices at A′(0, 0), B′(0, 10), and C′(15, 0). The origin is the center of dilation. What is the scale factor of the dilation?

Hoodmemes:

Ok so a scale factor k = 10/4 = 15/6 = (blank) hope this helps